OpenStudy (anonymous):

Ming & Houng each improved their yards by planting hostas & heraniums. They bought their supplies from the same store. Ming spent $102 on 6 hostas and 6 geraniums. Huong spent $93 on 6 hostas and 5 geraniums. Find the cost of one hosta and the cost of one geranium.

OpenStudy (anonymous):

let cost of one hosta is x and the cost of one geranium is y. and Ming spent $102 on 6 hostas and 6 geraniums so we can write this statement as equation 6x+6y =120 also Huong spent $93 on 6 hostas and 5 geraniums so 6x+5y= 93 now by salving thease two equation you will get x and y.

OpenStudy (anonymous):

what happened @joseemtez are u get it or not?

OpenStudy (anonymous):

I got y (geranium) its 6 right ? But i cant find x (hosta) im stuck

OpenStudy (anonymous):

@gyanu

OpenStudy (anonymous):

my apology for i write 120 in place of 102 correct solution is here let solve 6x+6y =102 .................(1) 6x+5y= 93 ...................(2) subtract (2) from (1) we get y=102-93=9 so y= 9 put value of y in (1) we get 6x+6*9=120 6x+54=102 so 6x=48 x=8

OpenStudy (anonymous):

Alright . Thanks